OpenStudy (anonymous):

HELP PLEASE. ALG 1 Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6.

OpenStudy (anonymous):

x intercept = 2 y intercept =-3

OpenStudy (anonymous):

How did you get that? @zordoloom

OpenStudy (anonymous):

Plug a zero into x if you want to find y. So, 3(0)-2y=-6

OpenStudy (anonymous):

And is it x-intercept: (2, 0), y-intercept: (0, −3) or x-intercept: (0, 2), y-intercept: (0, −3)

OpenStudy (anonymous):

(2,0) and (0,-3)

OpenStudy (anonymous):

@Haleyyy x-intercept itself means y-coordinate=0 so x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (anonymous):

I just don't understand how to do this stuff. My teacher taught it to me I just dont understand it and now i'm doing the hw and dont know how to work it out. Last Q: Choose the correct x-intercept and y-intercept of the equation x + 2y = 4.

OpenStudy (anonymous):

@Haleyyy then let make it to understand u:)

OpenStudy (anonymous):

@zordoloom don't post answers plz:)

OpenStudy (anonymous):

Okay step by step: Given: x + 2y = 4. To find the x intercept, replace the y with a zero: x+2(0)=4. x+0=4.....

OpenStudy (anonymous):

@zordoloom she don't know anything about intercept so wait plz:)

OpenStudy (anonymous):

okay. I'll let you take over.

OpenStudy (anonymous):

@Haleyyy do u know what is actually intercept and a line means??????

OpenStudy (anonymous):

yes, i just need to know how to work out the problem thats all.

OpenStudy (anonymous):

well then what is x-intercept means???????

OpenStudy (anonymous):

you're not making any sense. just take me step by step please like how @ zordoloom was :-)

OpenStudy (anonymous):

what do i do when i start with \[\chi +2y =4\]

OpenStudy (anonymous):

@Haleyyy line is just a sequence of points which means all points on it will satisfy and x-intercept means at which it cuts x-axis means value of y=0 so to find a point on the line where y=0 it must satisfy equatin so putting y=0 we find x so here x+2*(0)=4 x+0=4 x=4 so co-ordinates is(4,0) now try for y- intercept:)

OpenStudy (anonymous):

i figured it out! thx